OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y < -x + 6?

OpenStudy (anonymous):

options????

OpenStudy (anonymous):

(-3, 0) (-1, 0) (4, 3) (3, -2)

OpenStudy (anonymous):

Plug in the coordinates and see. I'll do one as an example... if we wish to test \((-3,0)\), we substitute \(x=-3\) and \(y=0\) into the inequality and test to see if it's true:$$y<-x+6\\0<-(-3)+6\\0<3+7\\0<9$$... which means \((-3,0)\) satisfies the inequality and is NOT the answer.

OpenStudy (anonymous):

BTW I meant 3 + 6 not 3 + 7

OpenStudy (rosedewittbukater):

Plug in all of the options as (x, y) and see which ones are true. The one that makes the equation not true is the answer. 0 < 3 + 6, =0<9 is true so that is not the answer. zero is less than nine 0 < 1 + 6 =0<7 is true so that is also not the answer. zero is less than seven 3 < -4 + 6 =3 < 2 is not true, so (4, 3) is the answer*** three is not less than two -2 < -3 + 6 = -2 < 3 is true so it is not the answer. negative two is less than 3